OpenStudy (anonymous):

A coin has a diameter of 26.5 mm. What is the approximate area of one side of the coin . A. 551mm squared B. 2205 mm squared C. 83 mm squared D.166 mm squared

OpenStudy (zale101):

Area of a circle when knowing the diameter A = (π/4) × d^2 A = (π/4) × 26.5^2

OpenStudy (zale101):

can u solve that?